Question 11)

A population of values has a normal distribution with μ=104.6 and σ=99.7. You intend to draw a random sample of size n=229.

Find P66, which is the mean separating the bottom 66% means from the top 34% means.
P66 (for sample means) =

Enter your answers as numbers accurate to 1 decimal place. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Given that,

mean = = 104.6

standard deviation = = 99.7

n = 229

= 104.6

= / n = 99.7 / 229=6.5884

Using standard normal table,

P(Z < z) = 66%

= P(Z < z) = 0.66  

= P(Z <0.412 ) = 0.66

z = 0.412

Using z-score formula  

= z * +

= 0.412*6.5884+104.6

= 107.3